OpenStudy (anonymous):

San FRancisco is known for its fine chocolates. If a deluxe box of fancy chocolates weighs 1 pounds 8 ounces and contains 32 chocolates piece, about how much does each piece weigh?

OpenStudy (anonymous):

a- 1 pound 8 ounces x 32 = n b- 1 pound + 8 ounces = n c- ( 6 ounces x 1) +8 = n d- 32/ 1 pound 8 ounces =n